OpenStudy (anonymous):

I know the answer is 1/7 but I don't understand whay can someone help. Suppose you select 2 letters at random from the word compute. Find each probability. P(2 vowels)

OpenStudy (anonymous):

\[P(2V)=\frac{\left(\begin{matrix}3 \\ 2\end{matrix}\right)}{\left(\begin{matrix}7 \\ 2\end{matrix}\right)}=\frac{1}{7}\] There are three vowels total, out of which you have to pick TWO vowels, and there are seven total letters, out of which you have to pick TWO letters...